# Dark Energy predictions from GREA: Background and linear perturbation theory

### v1

@@ Abstract
General Relativistic Entropic Acceleration (GREA) theory Espinosa-Portalés and García-Bellido (2021) gives a covariant formalism for out of equilibrium dynamics in the context of general relativity. A consequence of GREA is the explicit breaking of the time reversal invariance when there is entropy production. This drives an entropic force that behaves effectively like bulk viscosity with a negative effective pressure. The natural consequence of this formalism is that the source for space-time curvature is the Helmholtz free energy, $F=U-TS$, and not just matter and radiation García-Bellido (2023, 2024).
@@
In this paper, we perform a detailed calculation of the background and linear matter perturbations with the assumption that the entropy growth responsible for the accelerated expansion comes from the homogeneous cosmological horizon. We show that, in this case, the dynamics can be described in terms of just one parameter, $\alpha$, the ratio of the spatial curvature to the horizon distance, which gives the size of the cosmological horizon and the rate of growth of the horizon entropy. A given value of $\alpha$ defines a unique history of the universe, with a dynamics that differs from $\Lambda$CDM and gives and explicit prediction for all background observables, like the values of the matter and dark energy content of the universe, the rate of expansion today, the coasting point, the value of the equation of state of dark energy and its derivative, the angular diameter and luminosity distances, etc.
@@ II.1 The causal cosmological horizon
We can then solve the dynamical Friedmann equations for the accelerated universe in the context of an entropic force arising from the causal horizon. We will assume here a concrete scenario García-Bellido and Espinosa-Portalés (2021) in which the causal horizon coincides with the boundary of a bubble wall separating our flat FLRW universe from an open empty space, therefore matching conditions require that both the scale factor and its derivative should be continuous at that hypersurface, i.e. $3H_{\rm in}^{2}=\kappa\,\rho(a_{\rm in})$ and $H_{\rm out}^{2}=-k/a_{\rm in}^{2}$, which determines $\sqrt{-k}=a_{0}H(0)$ today. We write the Hamiltonian constraint in conformal time (where primes denote derivatives w.r.t. $\tau=\sqrt{-k}\,\eta=a_{0}H_{0}\eta$) as  [^1] Note that there is a difference w.r.t. Ref. García-Bellido and Espinosa-Portalés (2021), and we can now set $\Omega_{\rm K}=0$ inside the causal patch.
@@
We can then fix $\eta_{0}$ and determine the ratio between the causal horizon today and the curvature scale,
$\alpha\equiv\frac{a_{0}\,\eta_{0}}{a_{0}/\sqrt{-k}}=\sqrt{-k}\,\eta_{0}\,.$ (15)
@@
$d_{H}(a)=\frac{2}{\sqrt{\Omega}_{\rm M}}\,a^{3/2}\cdot{}_{2}F_{1}\left[\frac{1}{2},\,\frac{1}{6},\,\frac{7}{6},\,\frac{\Omega_{\rm M}-1}{\Omega_{\rm M}}\,a^{3}\right]\,,$ (16)
@@
 1.000  0.310  0.690   0.678  -1.000   -0.000  0.645
Table 1: The GREA values of the DM and DE content, the rate of expansion today, $H(0)=100\,h_{0}$ km/s/Mpc, the effective EOS parameters ($w_{0},\,w_{a}$) and the redshift $z_{c}$ of the corresponding coasting point, for different values of the parameter $\alpha$. We compare those values with the corresponding ones for $\Lambda$CDM (bottom row).
It is important to emphasize that GREA has exactly the same number of free parameters as $\Lambda$CDM, i.e. ($\alpha,\,\Omega_{\rm M},\,H_{0}$) versus ($\Omega_{\Lambda},\,\Omega_{\rm M},\,H_{0}$), and nevertheless has very different dynamics, both at the present times and in the far future. Moreover, while in $\Lambda$CDM there is a coincidence (“why now”) problem, i.e. why is $\Omega_{\Lambda}\simeq\Omega_{\rm M}$ precisely today?, in the case of GREA the growth of the cosmic horizon during the matter era is responsible for the transition to a dominance of the entropic repulsive forces associated with the acceleration of the universe. The moment at which this happens depends on the actual value of the spatial boundary today in units of the causal horizon, as can be seen in Fig. 2a, where we show the coasting points for the different realizations of GREA at different values of $\alpha$. Note that different GREA scenarios populate the parameter space of the cosmological background in a way that can be easily distinguished from $\Lambda$CDM, see Fig. 3.
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Furthermore, in GREA the value of the rate of expansion today is not fixed at the CMB, like in the case of $\Lambda$CDM. Here, depending on $\alpha$, the present rate of expansion, $H(0)$, could be larger or smaller than the $\Lambda$CDM value derived from the CMB, see Fig. 4a. In fact, for certain values of the spatial curvature one can resolve in a natural way the so-called Hubble tension Di Valentino et al. (2021).
Finally, while in $\Lambda$CDM, or its variant $w_{0}w_{a}\Lambda$CDM, the value of the equation of state of Dark Energy is given by $w(a)=w_{0}+w_{a}(1-a)$, as an expansion around the present value, in GREA we have the whole function $w(a)$, as can be seen in Fig. 2b. The evolution of the scale factor in GREA allows one to compute the effective equation of state of dark energy,
@@
In fact, in GREA theory there is no freedom to chose ($\Omega_{\rm M},\,H(0),\,w_{0},\,w_{a}$) independently, as it happens in $w_{0}w_{a}\Lambda$CDM. The actual values of $(w_{0},\,w_{a})$ are precisely given by the theory, just like the present rate of expansion, $H(0)=100\,h_{0}$ km/s/Mpc, once you fix the ratio $\alpha$, see Fig. 3 for some examples. As a consequence, if the cosmological data favours values that are far away from $\Lambda$CDM $(w_{0}=-1,\,w_{a}=0)$, then we can test the whole theory by comparing those values with the corresponding predictions for both $H(0)$ and ($\Omega_{\rm M},\,\Omega_{\Lambda}$). This can be seen in Table 1 and Fig. 5.
@@ IV.2 The BAO scale and the $S_{8}$ tension
Another cosmological observable which is derived from weak lensing is the magnitude of the fluctuations today on a scale of 8 Mpc/h. For $\Lambda$CDM and GREA we have different predictions and one can write the ratio as a function of $\alpha$, as shown in Fig. 9. For those values of $\alpha$ that resolve the $H_{0}$ tension, predicting today $H(0)=73\pm 1$ km/s/Mpc Di Valentino et al. (2021), the $\sigma_{8}$ value predicted by GREA is around 4% lower than that of $\Lambda$CDM, resolving also the $\sigma_{8}$ tension Di Valentino et al. (2021).
@@ IV.3 The ISW effect
The entropic acceleration due to the causal horizon starts to be important only recently, thanks to the $\sinh(2\tau)$ dependence in Eq. (13), resolving the coincidence problem: It is just the fact that the universe is old and big that GREA starts to dominate the expansion of the universe. In the past, such a term was exponentially negligible. Moreover, the effective nature of dark energy generates a time evolution that could be distinguished from that of the Standard Cosmological Model.

### v2

@@ Abstract
General Relativistic Entropic Acceleration (GREA) theory Espinosa-Portalés and García-Bellido (2021) gives a covariant formalism for out of equilibrium dynamics in the context of general relativity. A consequence of GREA is the explicit breaking of the time reversal invariance when there is entropy production. This drives an entropic force that behaves effectively like bulk viscosity with a negative effective pressure. The natural consequence of this formalism is that the source for space-time curvature is the Helmholtz free energy, $F=U-TS$, and not just matter and radiation García-Bellido (2024, 2023).
@@
In this paper, we perform a detailed calculation of the background and linear matter perturbations with the assumption that the entropy growth responsible for the accelerated expansion comes from the homogeneous cosmological horizon. We show that, in this case, the dynamics can be described in terms of just one parameter, $\alpha$, the ratio of the spatial curvature to the horizon distance, which gives the size of the cosmological horizon and the rate of growth of the horizon entropy. A given value of $\alpha$ defines a unique history of the universe, with a dynamics that differs from $\Lambda$CDM and gives an explicit prediction for all background observables, like the values of the matter and dark energy content of the universe, the rate of expansion today, the coasting point, the value of the equation of state of dark energy and its derivative, the angular diameter and luminosity distances, etc.
@@ II.1 The causal cosmological horizon
We can then solve the dynamical Friedmann equations for the accelerated universe in the context of an entropic force arising from the causal horizon. We will assume here a concrete scenario of open inflation García-Bellido and Espinosa-Portalés (2021) in which the causal horizon coincides with the boundary of a bubble wall separating our flat FLRW universe from an open empty space, therefore matching conditions require that both the scale factor and its derivative should be continuous at that hypersurface, i.e. $3H_{\rm in}^{2}=\kappa\,\rho(a_{\rm in})$ and $H_{\rm out}^{2}=-k/a_{\rm in}^{2}$, which determines $\sqrt{-k}=a_{0}H(a_{0})$ today. We write the Hamiltonian constraint in conformal time (where primes denote derivatives w.r.t. $\tau=\sqrt{-k}\,\eta=a_{0}H_{0}\eta$) as  [^1] Note that there is a difference w.r.t. Ref. García-Bellido and Espinosa-Portalés (2021), and we can now set $\Omega_{\rm K}=0$ inside the causal patch.
@@
We can then fix $\eta_{0}$ and determine the ratio between the causal horizon today and the curvature scale ($c=1$),
$\alpha\,H(a_{0})\,d_{H}(a_{0})\equiv\frac{a_{0}\,\eta_{0}}{a_{0}/\sqrt{-k}}=\sqrt{-k}\,\eta_{0}\,.$ (15)
@@
$H_{0}\,d_{H}(a)=\frac{2}{\sqrt{\Omega}_{\rm M}}\,a^{3/2}\cdot{}_{2}F_{1}\left[\frac{1}{2},\,\frac{1}{6},\,\frac{7}{6},\,\frac{\Omega_{\rm M}-1}{\Omega_{\rm M}}\,a^{3}\right]\,,$ (16)
@@
 0.310  0.690   0.678  -1.000   -0.000  0.645
Table 1: The GREA values of the DM and DE content, the rate of expansion today, $H(0)=100\,h_{0}$ km/s/Mpc, the effective EOS parameters ($w_{0},\,w_{a}$) and the redshift $z_{c}$ of the corresponding coasting point, for different values of the parameter $\alpha$. From top to bottom, the values of $\alpha$ correspond to the (red, orange, yellow, green, blue, purple) colors in the rest of the figures. We compare those values with the corresponding ones for $\Lambda$CDM (bottom row).
It is important to emphasize that GREA has exactly the same number of free parameters as $\Lambda$CDM, i.e. ($\alpha,\,\Omega_{\rm M},\,H_{0}$) versus ($\Omega_{\Lambda},\,\Omega_{\rm M},\,H_{0}$), and nevertheless has very different dynamics, both at the present times and in the far future. Moreover, while in $\Lambda$CDM there is a coincidence (“why now”) problem, i.e. why is $\Omega_{\Lambda}\simeq\Omega_{\rm M}$ precisely today?, in the case of GREA the growth of the cosmic horizon during the matter era is responsible for the transition to a dominance of the entropic repulsive forces associated with the acceleration of the universe. The moment at which this happens depends on the actual value of the spatial boundary today in units of the causal horizon, as can be seen in Fig. 2a, where we show the coasting points for the different realizations of GREA. In this scenario, the value of $\alpha$ depends on how far is our worldline from the bubble wall, as pictured in Fig. 1 of Ref.García-Bellido and Espinosa-Portalés (2021). If the bubble wall would have been further away, the onset of acceleration would have happened later for the same value of $\Omega_{\rm M}$ inside our universe. However, in an open universe the largest volume is near the bubble wall, so it is expected that most observers will lie far from the center of the inflated bubble and near its edges. Therefore the finetuning associated with the coincidence problem is strongly alleviated. Note also that different GREA scenarios populate the parameter space of the cosmological background in a way that can be easily distinguished from $\Lambda$CDM, see Fig. 3, with multiple panels, showing that of every 2D parameter space there is always a value of $\alpha$ that can accommodate a universe like that of $\Lambda$CDM, except for the value of $w_{a}$, which is always large and negative, in agreement with recent observations.
@@
Furthermore, in GREA the value of the rate of expansion today is not fixed at the CMB, like in the case of $\Lambda$CDM. Here, depending on $\alpha$, the present rate of expansion, $H(0)$, could be larger or smaller than the $\Lambda$CDM value derived from the CMB, see Fig. 4a and Table I. In fact, for certain values of the spatial curvature one can resolve in a natural way the so-called Hubble tension Di Valentino et al. (2021).
Finally, while in $\Lambda$CDM, or its variant $w_{0}w_{a}\Lambda$CDM, the value of the equation of state of Dark Energy is given by $w(a)=w_{0}+w_{a}(1-a)$, as an expansion around the present value, in GREA we have the whole function $w(a)$ to compare with observations, as can be seen in Fig. 2b. The evolution of the scale factor in GREA allows one to compute the effective equation of state of dark energy,
@@
In fact, in GREA theory there is no freedom to chose ($\Omega_{\rm M},\,H(0),\,w_{0},\,w_{a}$) independently, as it happens in $w_{0}w_{a}\Lambda$CDM. The actual values of $(w_{0},\,w_{a})$ are precisely given by the theory, just like the present rate of expansion, $H(0)=100\,h_{0}$ km/s/Mpc, once you fix the ratio $\alpha$, see Fig. 3 for some examples. As a consequence, if the cosmological data favours values that are far away from $\Lambda$CDM $(w_{0}=-1,\,w_{a}=0)$, then we can test the whole theory by comparing those values with the corresponding predictions for both $H(0)$ and ($\Omega_{\rm M},\,\Omega_{\Lambda}$), as can be seen in Table 1.
We note in Fig. 5 that the GREA scenario is not in tension with the recent DESI-BAO Adame et al. (2024) + DES-SNIa Abbott et al. (2024) data, although a proper analysis, including the background and linear perturbations of GREA, has to be performed with the full data set (CMB, SNIa, LSS, BAO, RSD, ISW, $H_{0}$, etc.). In fact, the amount of information that we can extract from present observations is larger than what can be inferred from their projection on to the plane ($w_{0},\,w_{a}$), and future data will be even more constraining at different redshifts.
@@ IV.2 The BAO scale and the $S_{8}$ tension
Another cosmological observable which is derived from weak lensing is the magnitude of the fluctuations today on a scale of 8 Mpc/h. For $\Lambda$CDM and GREA we have different predictions and one can write the ratio as a function of $\alpha$, as shown in Fig. 9. For those values of $\alpha$ that resolve the $H_{0}$ tension, predicting today $H(0)=73\pm 1$ km/s/Mpc Di Valentino et al. (2021), the $\sigma_{8}$ value predicted by GREA is around 4% lower than that of $\Lambda$CDM, resolving also the $S_{8}$ tension Di Valentino et al. (2021).
We should be cautious here, since in order to claim a resolution of the so-called $H_{0}$ and $S_{8}$ tensions within GREA requires a full CMB, weak lensing and Cepheid-TRGB analysis. In fact, recent claims of resolution of each of these tensions have been discussed recently and it could be that they are associated with unknown systematics yet to be understood.
@@ IV.3 The ISW effect
The entropic acceleration due to the causal horizon starts to be important only recently, thanks to the $\sinh(2\tau)$ dependence in Eq. (13), alleviating the coincidence problem. It is the fact that the universe is old and big that GREA started to dominate the expansion of the universe only recently. In the past, such a term was exponentially negligible, for ratios $\alpha\simeq 1$. One only has to live near the edge of the bubble, where most of the volume is. For larger horizon distances the acceleration would have started later for the same amount of matter.